Mountain Dental Services is a specialized dental practice whose only service is filling cavities. Mountain has recorded the following for the past nine months Month January February March April May June Jul August September Number of Cavities Filled 475 575 350 675 500 400 625 650 450 Total Cost $5,750 6,250 5,200 6,500 5,650 5,250 6,300 6,000 5,450 Required 1. Use the high-low method to estimate total fixed cost and variable cost per cavity filled. (Round your answers to 2 decimal places.) Fixed Cost Variable Cost per Unit 2. Using these estimates, calculate Mountains total cost for filling 550 cavities Estimated Total Cost
Answer #1

Variable cost per unit=[Total cost at highest level-Total cost at lowest level]/(Highest level-Lowest level)

=(6500-5200)/(675-350)

=$4 per unit

Total fixed cost=6500-(4*675)

=$3800

Total cost for 550 units=3800+(4*550)

=$6000.

NOTE:Total fixed cost and variable cost/unit do not change with change in units.